How much water should be added to 77 gallons of a 37% salt solution to get a 33% salt solution?
Answer by Onmyoji.S(32)

How much water should be added to 77 gallons of a 37% salt solution to get a 33% salt solution?
_________❤________________❤_________________❤______________
37%/77=33%/x
37/7700=33%/x
37/7700=33/100x
1/x=37/2541
x=2541/37
x approximately equal to 68.6
❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤
———————♥——————————//(◕‿‿◕)\\—————————♥———————